A loaded crane has a total mass of 225 kg. If a 312-N force acting at an angle of 30 degrees to the ground is applied, how much is done in moving the crane 20 feet? What is the weight of the loaded crane?


1
Expert's answer
2021-06-11T11:21:12-0400

Solution.

"m=225kg;"

"F=312N;"

"\theta=30^o;"

"s=20ft=6.096m;"

"W=Fscos\theta;"

"W=312N\sdot6.096m\sdot0.866=1647J;"

"P=mg;"

"P=225kg\sdot9.8N/kg=2205N;"

Answer: "W=1647J;" "P=2205N."